OpenStudy (anonymous):

Solve for x. Round your answer to 2 decimal places.

OpenStudy (anonymous):



OpenStudy (anonymous):

the bottom number is a 12

OpenStudy (anonymous):

I give medals :)

OpenStudy (anonymous):

So we know that cos is adj/hyp.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thus: \[\cos(51)=\frac{ 12 }{ x }\]

OpenStudy (anonymous):

51 degrees that is

OpenStudy (anonymous):

You can then switch the position of 'x' and cos(51). Giving you \[x = \frac{ 12 }{ \cos(51) }\]

OpenStudy (anonymous):

Plug that into your handy dandy calculator and you will get 19.07 (rounded)

OpenStudy (anonymous):

Hope this helps :)

OpenStudy (anonymous):

Thank-You so much!

OpenStudy (anonymous):

No Problem!